OpenStudy (anonymous):

Am I right? 4. If a positive integer n, divided by 5 has a remainder 2, which of the following must be true? I n is odd II n + 1 cannot be a prime number III (n + 2) divided by 7 has remainder 2 A. none B. I only C. I and II only D. II and III only E. I, II and III

OpenStudy (anonymous):

I got none, or A.

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

ya right a is correct

OpenStudy (anonymous):

5. A solid cube of side 6 is first painted pink and then cut into smaller cubes of side 2. How many of the smaller cubes have paint on exactly 2 sides? I got 8, right?

OpenStudy (hihi67):

Yes, that is correct.

OpenStudy (anonymous):

ty!

OpenStudy (hihi67):

No problem.

OpenStudy (anonymous):

6. Line l contains the points (3,1) and (4,4). If line m is a different line, parallel to line l in the same coordinate plane, which of the following could be the equation of line m? A. y = 3x - 8 B. y = 1/3x - 3 C. y = -3x - 8 D. y = 3x + 1 E. y = -8x + 3 I got D, why can't be A though?

OpenStudy (mathstudent55):

5. A solid cube of side 6 is first painted pink and then cut into smaller cubes of side 2. How many of the smaller cubes have paint on exactly 2 sides? I got 8, right?  Look in the figure above. Each edge has a middle 2 x 2 cube with 2 faces painted pink. When you take into account the top face and the bottom face, their 8 edges have 8 cubes with 2 faces painted pink. Then each of the 4 vertical edges also has a cube painted pink on two faces, so the total is 12, not 8.